OpenStudy (anonymous):

A clock strikes once at 1:00, twice at 2:00, three times at 3:00 and so forth. How many times does it strike over a 12 hour period?

OpenStudy (anonymous):

so we are finding the sum?

OpenStudy (anonymous):

1+2+3+4+5+6+7+8+9+10+11+12?

sam (.sam.):

Alright we can set that up like 1,2,3,4,5...12

sam (.sam.):

You can do that too but that's not the way of answering it

OpenStudy (anonymous):

DaWizjr math 20-1?

sam (.sam.):

Since we need the sum of those numbers, the common difference here is 1, \[\large S_n=\frac{n}{2}[2a_1+(n-1)d] \\ \\ \large S_{12}=\frac{12}{2}[2(1)+(12-1)1]=?\]

sam (.sam.):

Try that out :)

OpenStudy (anonymous):

N=12 a1 or t1=1:00 d=?

sam (.sam.):

d is 1, because 2-1=1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

12=12/2[2(1)+(12-1)1]

OpenStudy (anonymous):

guessing t1=1?

sam (.sam.):

noo \(S_{12}\) is just a statement, \[\frac{12}{2}[2(1)+(12-1)1]=?\]

OpenStudy (anonymous):

Yea

sam (.sam.):

\(T_1=a_1=1\)

OpenStudy (anonymous):

I got 78

sam (.sam.):

Correct

OpenStudy (anonymous):

Nice